Blog from August, 2011

8-28-2011 SDSC Relocation

The following servers/services should be operating nominally:

  • Everything

Delivery of the switch hardware we need was delayed. It did not arrive until 11:30am. As a result, there has been a corresponding slip in our schedule.

List of affected Virtual Machines (VMs)

A list of affected Virtual Machines (VMs) can be found here

Project Information


Intermittent network interruptions while the network is upgraded this morning.
Oracle databases and Oracle database servers unavailable during NetApp move this afternoon.
victory and constellation oracle servers have been relocated.


Intermittent network interruptions while the network is upgraded.
While the maunaloa storage system is moved, the following data stores will be unavailable.

  • CellImageLibrary
  • HarvardData
  • Image Server "scratch" space


Intermittent network interruptions while the network is upgraded.

  • NIF1, NIF2, NIF4 and nif-crawler servers have been patched, updated and relocated.


Intermittent network interruptions while the network is upgraded.
Aproximately 30 minute outage while bare metal server is relocated.

Work to be Done

switch hardware upgrade

Delivery of the switch hardware we need was delayed. It is due by noon today, via FedEx.

"Bare Metal" servers

Servers that are not virtualized will be moved this morning while we are installing the new switch hardware. This will impact:

  • braininfo
  • the Oracle 3-node RAC system and databases hosted there.
  • maunaloa data storage
    • the SVN data repository
    • the CVS data repository

VM migration status

We are hoping the switches will arrive early enough to allow us to migrate the VMs over 10Gb, suspend them over 10Gb while we relocate the NetApp, and then resume them.

Oracle RAC move

Daniel Wei will be assisting us with the move of this equipment.

NetApp move

The NetApp will be one of the last pieces of equipment to be relocated. During this time, all VMs that use shared NetApp storage to facilitate disaster recovery will be suspended. CAMERA Oracle databases will be unavailable. WBC data stored on the NetApp will also be unavailable.

Work completed

VM migration

Four of our VM hosts have now been moved to the new location and a number of VMs have been migrated to these machines.

General Improvements

  • BIRN Portal and GAMA servers have been virtualized
  • NIF server system software updated

NIF status

  • NIF1, NIF2, NIF4, NIF5 have all been moved to the new location. Patches have been applied and network has been reconfigured.  Network has improved with adding redudent/failover links for all of the machines for both management and public networks.  Management network upgraded to 1GB.
  • Currently I have NIF's website pointing to a different webserver so that when the production nodes go down, we will still have a status page up for NIF.

We have upgraded power and all the necessary network drops in the new location. Progress is being impeded by the lack of redundant 10Gb layer 3 switches for the new location. We've done everything we can think of to expedite the procurement of this equipment, including requesting Saturday morning delivery via FedEx, if necessary.

In addition, we have moved the first VM host to the new location and plan to start migrating VMs to it tomorrow. Unfortunately, without the switches, the process of live migrating the VMs will take much longer than originally planned for.


What is Going On

The Oracle RAC at SDSC did not come up properly after the power outage yesterday.

System Affected


Project(s) - Software/Application(s) Affected

CCDB - production iRODS (iCAT database is offline)

Steps Being Taken to Rectify the Situation

Daniel Wei has been contacted (left message) and his assistance has been requested. We are waiting for him to get in touch with us.

ETA (if known)